SOLVE : cos[sin^(-1) (2/5) + cos^(-1) (x)] = 0 plz give stepwise solution

SOLVE :  cos[sin^(-1) (2/5) + cos^(-1) (x)] = 0
plz give stepwise solution

Grade:12

1 Answers

moumi roy
7 years ago
given:cos[sin^(-1) (2/5) + cos^(-1) (x)] = 0
cos[sin^(-1) (2/5) + cos^(-1) (x)]=cos[(pi)/2]
sin^(-1) (2/5) + cos^(-1) (x)=(pi)/2
so,x=(2/5)
as sin^(-1) (A) + cos^(-1) (A)=(pi)/2